### v4

# AICO: Feature Significance Tests for Supervised Learning

Abstract
Machine learning has become a central tool across scientific, industrial, and policy domains. Algorithms now identify chemical properties, forecast disease risk, screen borrowers, and guide public interventions. Yet this predictive power often comes at the cost of transparency: we rarely know which input features truly drive a model’s predictions. Without such understanding, researchers cannot draw reliable scientific conclusions, practitioners cannot ensure fairness or accountability, and policy makers cannot trust or govern model-based decisions. Despite its importance, existing tools for assessing feature influence are limited—most lack statistical guarantees, and many require costly retraining or surrogate modeling, making them impractical for large modern models. We introduce AICO, a broadly applicable framework that turns model interpretability into an efficient statistical exercise. AICO asks, for any trained regression or classification model, whether each feature genuinely improves model performance. It does so by masking the feature’s information and measuring the resulting change in performance. The method delivers exact, finite-sample inference—exact feature $p$-values and confidence intervals—without any retraining, surrogate modeling, or distributional assumptions, making it feasible for today’s large-scale algorithms. In both controlled experiments and real applications—from credit scoring to mortgage-behavior prediction—AICO consistently pinpoints the variables that drive model behavior, providing a fast and reliable path toward transparent and trustworthy machine learning.
1 Introduction
Machine learning has become integral to modern science, industry, and public policy. Algorithms now predict molecular activity, identify disease markers, forecast financial defaults, and guide the allocation of social resources. Yet even as these systems achieve extraordinary predictive accuracy, their inner workings remain largely opaque. We often do not know which input features truly shape a model’s predictions, which merely correlate by chance, or which may encode unwanted biases. This opacity limits scientific understanding, hinders regulatory oversight, and undermines public trust in data-driven decision making. Despite extensive progress in explainable artificial intelligence, most tools for measuring feature importance lack statistical guarantees or depend on computationally costly retraining or surrogate modeling procedures, restricting their use in today’s large, complex models. As a result, reconciling interpretability with scalable inference remains an open challenge for modern machine-learning practice.
We introduce AICO, a tractable method for feature significance testing in any supervised learning model, whether for regression or classification. AICO transforms the question of interpretability into one of formal statistical inference: it tests whether a feature contributes meaningfully to a model’s predictive or decision-making performance, using only the model’s fitted predictions and observed outcomes. AICO constructs test statistics based on the change in model performance when a feature is masked. It is model- and distribution-agnostic: it makes no assumptions about the model class, data distribution, or estimand (for example, conditional mean or quantile). The method does not rely on retraining or surrogate models, making it well suited for analyzing complex, high-dimensional algorithms that are costly to train. It bridges a long-standing gap between machine-learning practice and classical statistical inference, offering a tractable route to transparent, trustworthy prediction in today’s large-scale applications.
The AICO feature effect—the change in model performance due to masking—quantifies a feature’s importance for each individual sample. For a null feature, these effects form a distribution across test samples whose median is non-positive. A uniformly most powerful randomized sign test achieves finite-sample exactness: its size (type-I error probability) exactly matches the nominal level (e.g., $5\%$). The procedure produces exact, randomized feature $p$-values whose conditional distribution, given the data, is uniform on an interval that can be reported in practice. It is simple, robust, and computationally efficient, relying on minimal assumptions and remaining valid under broad conditions. These properties enable valid inference on a model’s features without relying on large-sample approximations or parametric assumptions.
The median feature effect being tested provides a natural, population-level measure of feature importance. We develop randomized, exact-coverage confidence intervals for this median that, together with the $p$-values, enable principled inference on a feature’s influence. The availability of confidence intervals distinguishes the median from widely used algorithmic measures of feature importance (see Section 1.1). The median also has several desirable properties that depend on the choice of masking value. It can be robust to correlation among features—a pervasive challenge in practical data settings—and, in linear regression, it is proportional to the square of the feature’s regression coefficient, the canonical measure of importance in the linear context. Finally, we provide conditions under which a feature’s median equal to zero is equivalent to conditional independence of that feature and the response given the remaining covariates.
Extensive experiments on synthetic “known-truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings. We compare AICO with powerful, model-agnostic, and non-asymptotic tests by ?, ?, and ?, as well as with the model-X knockoff procedures of ? and ? for variable selection under false discovery control. Our test achieves comparable or higher statistical power than these alternatives, with minimal false discoveries, while reducing computational cost by several orders of magnitude. We further show that the AICO feature-importance measure is robust to noise in the data, yielding stable feature rankings. This stands in sharp contrast to widely used algorithmic importance measures such as SHAP (?), which are highly sensitive to data perturbations and can produce unstable rankings that hinder reliable inference about variable importance. These empirical findings reinforce recent theoretical discussions of the limitations of algorithmic importance measures in ?, ?, and related work.
Empirical analyses demonstrate the ability of our framework to inform high-stakes decisions with legal, economic, and regulatory implications. In a credit scoring task using ensembles of gradient-boosting models, AICO identifies and ranks several statistically significant features describing past payment behavior and social factors. Notably, the analysis reveals that gender, marital status, and age of card holders—protected characteristics under U.S. lending laws—are statistically significant predictors of card default, even after controlling for past payment behavior. In a mortgage state–transition prediction task based on a recently proposed set–sequence model with long-convolution architectures (?) and trained on extensive panel data, AICO detects time-varying macroeconomic variables such as mortgage rates, as well as features summarizing past state transitions, as significant predictors of borrower behavior.
The remainder of this introduction discusses the related literature. Section 2 introduces the notion of feature effects, and Section 3 develops the testing procedure. Section 4 analyzes the median feature effect as an importance measure dual to the test. Section 5 examines the performance of the test and importance measure in synthetic regression and classification settings, while Section 6 applies the method to two real-world data sets. Section 7 concludes. Proofs of the theoretical results are provided in an Appendix. An open-source Python package implementing the testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
1.1 Related literature
There is a large body of work developing algorithmic feature importance measures; see ? for an overview and ? for a unifying framework for interpretability methods. Popular examples include measures based on Shapley values (e.g. SHAP, see ? and many others), measures based on local approximations (e.g. LIME, see ?), and measures obtained by permutation methods (e.g. ?, ?). These measures provide valuable insight into the role of features but unlike the feature importance measures delivered by our hypothesis testing approach, they typically do not offer statistical guarantees or $p$-values. We demonstrate that algorithmic measures can be highly susceptible to noise in the data so the lack of confidence intervals is concerning. Moreover, algorithmic measures are often biased in the presence of feature correlation (?, ?). Our approach can deliver importance measures that are robust to feature correlation. Finally, we demonstrate that our method achieves orders-of-magnitude improvements in computational efficiency over widely used algorithmic importance approaches, such as SHAP and permutation scores.
The statistical and econometric literature has long studied significance testing for variables in nonparametric regression models. ?, ?, ?, ?, ?, and others employ kernel-based methods to construct asymptotic testing procedures. ?, ?, ?, and others develop gradient-based asymptotic tests using neural networks. ? develop permutation-based asymptotic tests for random forests; see also ? for an alternative approach using $U$-statistics. In contrast, the AICO tests are model-agnostic and non-asymptotic, deliver exact $p$-values and confidence intervals, and accommodate a broad range of data structures.
Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ?, and others develop non-asymptotic tests for conditional independence in the model-X knockoff framework of ?. Conditional independence testing is inherently difficult: ? show that no such test can be uniformly valid against arbitrary alternatives. We show that, for an appropriate choice of the feature masking value, the AICO null is generally weaker than the conditional-independence null studied in these papers, with equivalence holding under additional conditions. ? suggest model-agnostic tests based on the drop in model predictive accuracy when features are removed from the model (LOCO: leave-out-covariates; see also ? and ?). In contrast, AICO tests the change in model performance due to masking a feature.
The computational demands of the aforementioned model-agnostic testing procedures are a major limiting factor. The LOCO test requires retraining the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train. The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, a task that can be as challenging and computationally expensive as fitting the predictive model itself. Indeed, ? show that no algorithm can efficiently compute nontrivial knockoffs for arbitrary feature distributions. In contrast, AICO requires neither model retraining nor auxiliary modeling, delivering an unprecedented degree of computational efficiency. Our numerical experiments demonstrate that these drastic gains in computational speed do not come at the expense of statistical power or inferential accuracy.
2 Feature effect
Consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$. Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a feature vector taking values in a set $\mathcal{X}$, and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$. For example, in a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$. We make no assumptions on the distribution of the $Z_{i}$’s.
We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, used for training and testing, respectively. Such data splitting is standard in supervised learning and, in our setting, facilitates post-selection inference.
We also take as given a model (algorithm) $f:\mathcal{X}\to\mathcal{M}$ trained on $\{Z_{i}\}_{i\in I_{1}}$, which is to be tested. We make no assumptions on the form of the estimand; $f$ may represent the conditional mean, a quantile, or another statistic of $Y_{i}$. We likewise do not impose assumptions on the training criterion (e.g., $L_{2}$ error) or on the optimization approach (e.g., empirical risk minimization) used to fit $f$.
We aim to test the statistical significance of a feature $\ell\in[d]$ under the fitted model $f$. (Unless noted otherwise, $\ell$ is fixed.) When $f$ is linear, classical asymptotic procedures such as the $t$-test are available. We develop instead a simple, non-asymptotic, and model-agnostic testing procedure that does not rely on a specific model structure and remains valid for models $f$ of essentially any form.
We consider the impact of a feature’s variability on model performance across the test set $I_{2}$. Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$ that relates model output to response. For a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ relative to $y$, with higher scores indicating better performance. A natural choice for $S$ is the negative of the loss function used to train $f$. Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in regression, the negative pinball loss in quantile regression, or the negative (binary or multi-class) cross-entropy loss in classification. With these specifications, $S$ measures predictive accuracy. Other choices of $S$ can quantify model performance in downstream screening or decisioning applications of $f$ (e.g., revenue or cost), such as the financial loss caused by a poor decision due to misclassification. Finally, one may specify a function $S$ that does not depend on the second argument $y$ to measure the behavior of the model output itself (for example, its sensitivity to the inputs).
We evaluate the effect of a feature’s variability on the model score across the test set. For each $i\in I_{2}$, we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$, serving as a reference, and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$. In the baseline, the value of the feature to be tested is masked with a reference value, the choice of which is discussed further below. The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can leverage the actual sample-$i$ value of the feature being tested. We measure the feature’s effect by the corresponding improvement in model score.
Definition 2.1 (Feature Effect).
For a given score function $S$, baseline feature value $\underline{X}_{i}$, and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is defined as the score difference
$\Delta_{i}=S(f(\overline{X}_{i}),Y_{i})-S(f(\underline{X}_{i}),Y_{i}).$
Masking feature values to measure a feature’s effect offers significant advantages over alternatives such as random shuffling, random perturbation or dropping the feature from the vector. Random procedures introduce additional variance into the measurement, while dropping features requires training sub-models, which drastically reduces computational efficiency. Our masking approach avoids both issues, yielding stable and computationally efficient estimates of feature importance.
2.1 Baseline and comparison features
Along with the selection of the score function $S$, the choice of baseline and comparison feature values allows users to tailor the feature effect to the model, data structure, and testing objectives at hand. We provide a conceptual framework and several examples. Throughout, fix a feature $\ell\in[d]$ to be tested; our notation usually suppresses the dependence on $\ell$.
We begin by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for each test sample $i\in I_{2}$. In a second step, we aggregate the samples in $\underline{I}$ to obtain $\underline{X}_{i}$. Possible aggregation rules include the mean, a quantile, or a mode of $\{X_{k}:k\in\underline{I}\}$.
Example 2.1.
Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$, where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$. Then $f(\underline{X}_{i})$ is constant across the test set. For $\ell=1$, take $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$. Here, the feature’s actual sample value $X_{i}^{1}$ is added back in while all other components remain fixed—an operation that gives rise to the name AICO (Add-In COvariates). The predicted response $f(\overline{X}_{i})$ therefore leverages the true variation in feature 1, and the feature effect $\Delta_{i}$ represents the resulting improvement in model score.
An alternative choice is $\underline{X}_{i}=(v^{1},X_{i}^{2},\ldots,X_{i}^{d})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d})$, again with $\ell=1$. Here, the predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, as it incorporates variation in the remaining features $X_{i}^{2},\ldots,X_{i}^{d}$. Adding the true value $X_{i}^{1}$ back into this context performs the same AICO operation, now yielding a conditional feature effect $\Delta_{i}$ that measures the incremental contribution of feature 1 given the others.
Choices other than the mean may also be appropriate for $v^{j}$. For instance, when $X_{i}^{j}\in\{0,1\}$ encodes a categorical variable, the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ may not represent a valid feature value. Sensible alternatives include the mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the complementary value $1-X_{i}^{j}$.
The set $\underline{I}$ need not equal the entire training set $I_{1}$. We can select specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then choosing the samples from $I_{1}$ that lie in or outside a neighborhood of $r$. The point $r$ can represent, for example, the (weighted) mean, mode, or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, of the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or of the feature values $\{X^{\ell}_{k}:k\in I_{1}\}$ for the feature being tested. Quantiles with $\lambda\in\{0,1\}$ are particularly interesting special cases. Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ a distance measure on $\mathcal{Y}$. We can then take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed (or surpasses) a given threshold. Other criteria can also be used, such as selecting samples for which $Y_{k}$ exceeds or falls short of $Y_{r}$. A similar approach applies when $r$ is a reference point in $\mathcal{M}$ or $\mathcal{X}$.
Example 2.2.
Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is the $\lambda=0$ quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$. In a binary soft-classification setting such as fraud detection or credit scoring, where $Y_{i}=0$ denotes the preferred (beneficial) outcome, this reference point corresponds to the training samples with minimum predicted risk. [^2] See ?. We can take $\underline{I}=\{\mbox{argmin}_{k\in I_{1}}f(Y_{k})\}$ as the set of these samples and define $\underline{X}_{i}$ and $\overline{X}_{i}$ as in Example 2.1. The resulting feature effect $\Delta_{i}$ represents the score delta relative to the minimum-risk baseline and measures how much excess risk a feature explains.
A few general principles guide the selection of reference points and the construction of $\underline{I}\subseteq I_{1}$. We aim to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, avoiding feature effects $\Delta_{i}$ that are biased toward zero and lead to significance tests with low power.333? discuss related issues in the context of image classification. At the same time, we seek to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Further, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to choose an appropriate reference point and aggregation rule for $\underline{I}$ (see Example 2.1). Alternatively, one may select a suitable baseline value directly from the set $\{X_{k}:k\in I_{1}\}$—namely, an element near the raw off-manifold value produced by the baseline selection procedure. [^4] We also allow for the selection and aggregation of alternative baselines for a given sample. A general approach is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
2.2 Multiple features, panel data and missing data
The discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The same framework naturally extends to subsets of $[d]$, enabling the joint testing of multiple features and the exploration of higher-order effects, including feature interactions.
Panel data settings—where collections of units are observed over time—are common in many domains. Appendix B discusses two simple extensions, conditioning and grouping, that adapt the construction of the feature effect to such data structures.
Real-world data sets often contain missing feature values. Appendix C describes how missing or imputed values can be handled when constructing $\Delta_{i}$.
3 Hypothesis testing
To construct a hypothesis test, we examine the empirical distribution of the feature effects $\Delta_{i}$ across the test samples $i\in I_{2}$, keeping the feature $\ell\in[d]$ fixed. If this distribution has a non-positive center, the feature is deemed insignificant; a large positive center provides evidence of significance. Several classical tests could be applied to assess such central tendencies, including $t$, sign, Wilcoxon, and permutation tests. We propose a simple sign test, which requires minimal assumptions on the data while retaining meaningful asymptotic relative efficiency compared with a $t$-test.
Assumption 3.1.
The feature effects $\{\Delta_{i}:i\in I_{2}\}$ are independent draws from a conditional distribution $\mathbb{P}_{M}$ given the training data $I_{1}$, with unique median $M\in\mathbb{R}$. The associated distribution function $F_{\Delta}$ is continuous. [^5] A sufficient condition for the uniqueness of $M$ is strict monotonicity of $F_{\Delta}$ in a neighborhood of $M$.
Let $B_{n,p}$ be a Binomial-$(n,p)$ variable and let $q_{\lambda}(B_{n,p})=\inf\{x\geq 0:P(B_{n,p}\leq x)\geq\lambda\}$ be its (lower) quantile at $\lambda\in(0,1)$. Let $N=|I_{2}|$ be the size of the test sample, and write $\Pi_{N}(n)=\sum_{i=0}^{n}\pi_{N}(i)$ for the CDF of $B_{N,1/2}$, where $\pi_{N}(i)={N\choose i}2^{-N}$ for $i=0,1,\ldots,N$.
Proposition 3.2.
Under Assumption 3.1, consider the hypotheses
$H_{0}:M\leq M_{0}\quad vs.\quad H_{1}:M>M_{0}$ (1)
for some $M_{0}\in\mathbb{R}$. A Uniformly Most Powerful test of size $\alpha\in(0,1)$ rejects $H_{0}$ with probability $\phi_{N}(n_{+}(M_{0}),\alpha)$ where $n_{+}(x)=\#\{i\in I_{2}:\Delta_{i}>x\}$ is the number of feature effect samples in the test set exceeding $x\in\mathbb{R}$ and
$\phi_{N}(n,\alpha)=1\{n>T_{N,\alpha}\}+\gamma_{N}(n,\alpha)1\{n=T_{N,\alpha}\}$ (2)
where $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ and
$\gamma_{N}(n,\alpha)=\frac{\Pi_{N}(n)-(1-\alpha)}{\pi_{N}(n)},\qquad n=0,1,\ldots,N.$ (3)
Input:
-0.5em
  • Model $f$ trained on $\{X_{i},Y_{i}\}_{i\in I_{1}}$
  • Baseline feature values $\{\underline{X}_{i}\}_{i\in I_{2}}$ for $\ell$
  • Comparison feature values $\{\overline{X}_{i}\}_{i\in I_{2}}$ for $\ell$
  • Response values $\{Y_{i}\}_{i\in I_{2}}$
  • Significance level $\alpha\in(0,1)$
  • Score function $S$
-0.5em Output:
-0.5em
  • Test conclusion (True = Significant at level $\alpha$, False = Insignificant at level $\alpha$)
-0.5em
Compute $\Delta_{i}=S(f(\overline{X}_{i}),Y_{i})-S(f(\underline{X}_{i}),Y_{i})$ for all $i\in I_{2}$ ;
Compute the test statistic $n_{+}=\#\{i\in I_{2}:\Delta_{i}>M_{0}\}$ ;
Compute the threshold $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ ;
if $n_{+}>T_{N,\alpha}$ then
 return True;
else if $n_{+}=T_{N,\alpha}$ then
   Generate $U\sim U(0,1)$ ;
   Compute the probability $\gamma_{N}(n_{+},\alpha)$ using (3) ;
 if $U\leq\gamma_{N}(n_{+},\alpha)$ then
    return True ;
return False
0.5em
Algorithm 1 AICO Test algorithm for feature $\ell\in[d]$.
For brevity, we often suppress the dependence of the test statistic $n_{+}$ on $M_{0}$. The procedure rejects $H_{0}$ with probability one if $n_{+}$—the number of feature effects in the test set exceeding $M_{0}$—is greater than the $(1-\alpha)$-quantile of the $\mathrm{Binomial}(N,1/2)$ distribution. If $n_{+}$ lies exactly on the boundary of the rejection region, the test randomizes. Relative to a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$, this randomization makes the procedure slightly more aggressive: it achieves the maximum possible power while guaranteeing exactness, in the sense that
$\mathbb{E}_{M_{0}}(\phi(n_{+},\alpha))=\alpha,\quad M_{0}\in\mathbb{R}.$ (4)
Algorithm 1 summarizes the implementation of the test. In our empirical applications in Sections 5 and 6, we use the value $M_{0}=0$ in (1).
In the context of our randomized test the notion of $p$-value is ambiguous. A conventional $p$-value could be defined as $\inf\{\alpha:\phi_{N}(n_{+},\alpha)=1\}$, the smallest level of significance $\alpha$ at which $H_{0}$ is rejected with probability one, which is $1-\Pi_{N}(n_{+}-1)$. (This is also a $p$-value for a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$.) However, this formulation ignores the randomness of the UMP test’s decision rule. We follow ? and consider a randomized $p$-value $p(n_{+})$ with conditional distribution function given by
$\mathbb{P}_{M_{0}}(p(n_{+})\leq\alpha\,|\,n_{+})=\phi_{N}(n_{+},\alpha),\quad\alpha\in(0,1).$
From (2), this function represents a uniform distribution on an interval $(a_{N}(n_{+}),b_{N}(n_{+}))$ of length $\pi_{N}(n_{+})$ where
$a_{N}(n)$ $=P(B_{N,1/2}>n)=1-\Pi_{N}(n),$
$b_{N}(n)$ $=P(B_{N,1/2}\geq n)=1-\Pi_{N}(n-1),$
the right endpoint being the conventional $p$-value. By iterated expectations and (4), we have
$\mathbb{P}_{M_{0}}(p(n_{+})\leq\alpha)=\mathbb{E}_{M_{0}}(\phi_{N}(n_{+},\alpha))=\alpha,\quad M_{0}\in\mathbb{R},$
meaning that $p(n_{+})\sim U(0,1)$. In this sense, the randomized $p$-value $p(n_{+})$ is exact, mirroring the exactness of the test’s decision rule (2). If $p(n_{+})$ is concentrated below a given level $\alpha$, or equivalently, if the conventional $p$-value $b_{N}(n_{+})\leq\alpha$, we reject the null. If $p(n_{+})$ is concentrated above $\alpha$, or equivalently, if $\alpha<a_{N}(n_{+})$, we retain the null. Else the evidence is equivocal; we thus randomize by drawing a sample $u$ from $U(a_{N}(n_{+}),b_{N}(n_{+}))$ and reject if $u\leq\alpha$. See Figure 1 for a visualization. In practice, it suffices to report the interval $(a_{N}(n_{+}),b_{N}(n_{+}))$ to fully characterize the $p$-value; there is often no need to force a decision.
[…]
Figure 1: Distribution of randomized $p$-value and decision intervals.
The power function $H_{N,\alpha}(M)$ of the test is given by
$H_{N,\alpha}(M)$ $=\mathbb{E}_{M}(\phi_{N}(n_{+}(M_{0}),\alpha))$
$=\mathbb{P}_{M}(n_{+}(M_{0})>T_{N,\alpha})+\gamma_{N}(T_{N,\alpha},\alpha)\mathbb{P}_{M}(n_{+}(M_{0})=T_{N,\alpha})$
$=P(B_{N,p}>T_{N,\alpha})+\gamma_{N}(T_{N,\alpha},\alpha)P(B_{N,p}=T_{N,\alpha})$ (5)
for values $M\in\mathbb{R}$, where $p=\mathbb{P}_{M}(\Delta_{i}>M_{0})$. The binomial probabilities in (3) are explicit. Since the test is uniformly most powerful, the function $H_{N,\alpha}$ is increasing for fixed $N$ and $\alpha$. We can invert that function to determine the size of the test set $N=|I_{2}|$ required to achieve a desired level of power. The parameter $p$ can be estimated using a pilot sample or other approach, as discussed by ?. Figure 2 plots the test sample size as a function of test power for several alternative values of $p$.
Figure 2: Test sample size NN vs. power $H_{N,\alpha}$ for $\alpha=5\%$ and several values of $p>1/2$.
The multiple testing problem can also be addressed in the randomized setting. A standard Bonferroni adjustment with test function $\phi_{N}(n,\alpha/d)$ and $p$-value being conditionally uniform on the interval $(da_{N}(n_{+})\wedge 1,db_{N}(n_{+})\wedge 1)$, where $d$ is the dimension of the feature vector, controls the family-wise error rate (FWER, the probability of having at least one false discovery), see ?. For a less conservative approach, one can also control the False Discovery Rate (?). ? develop a formula and an algorithm for the adjusted rejection probability. ? proposes an alternative Monte Carlo approach. ? provide a comparison of methods, finding that the Monte Carlo approach outperforms its alternatives in terms of test power.
4 Measuring feature importance
We seek measures of feature importance that complement the feature $p$-values delivered by our test. A local measure of feature importance is given by the feature effect $\Delta_{i}$, which captures the impact of a feature at the sample level. At the population level, a natural measure of feature importance is the median $M$ of the (conditional) feature-effect distribution $F_{\Delta}$. Under Assumption 3.1, a consistent estimator of $M$ is the empirical median
$\widehat{M}=\mathrm{Median}\{\Delta_{i}:i\in I_{2}\},$ (6)
which satisfies $\mathbb{P}(\widehat{M}<M)=\mathbb{P}(\widehat{M}>M)$ for odd $N=|I_{2}|$, with approximate equality for even $N$. Hence, (6) is (approximately) as likely to overestimate $M$ as to underestimate it. Moreover, the distribution of $\widehat{M}$ is continuous. The estimator $\widehat{M}$ is unbiased if $F_{\Delta}$ is symmetric about $M$; see Section 4.4 of ?.
We construct one-sided, randomized confidence intervals (CIs) for $M$ by inverting the test (2). Let $\{\Delta_{(i)}\}$ denote the ordered sequence of $\{\Delta_{i}\}$; we take $\Delta_{(0)}=$-$\infty$ and $\Delta_{(N+1)}=\infty$. Let $\mathbb{P}_{M}$ denote the sampling distribution of the $\{\Delta_{i}\}$ when $M$ is the true median.
Proposition 4.1.
Let $U$ be an independent $U(0,1)$ variable. Define a randomized CI by
$C_{N}(\alpha;U)=\begin{cases}[\Delta_{(N-T_{N,\alpha})},\infty),&1-\gamma_{N}(T_{N,\alpha},\alpha)\geq U\\ [\Delta_{(N-T_{N,\alpha}+1)},\infty),&\text{else}\end{cases}$ (7)
where $T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ is the boundary of the test’s rejection region and $\gamma_{N}(n,\alpha)$ is the rejection probability (3). Under Assumption 3.1, the CI (7) has exact coverage:
$\mathbb{P}_{M}(M\in C_{N}(\alpha;U))=1-\alpha,\quad\alpha\in(0,1).$ (8)
In analogy to the randomized $p$-value, the randomized CI (7) inherits the exactness of the test’s decision rule (2). If $0<\Delta_{(N-T_{N,\alpha})}$ we reject the null for $M_{0}=0$. If $\Delta_{(N-T_{N,\alpha}+1)}\leq 0$ we retain the null. Else we reject with probability $\gamma_{N}(T_{N,\alpha},\alpha)$ (that is, draw a sample $u$ from $U(0,1)$ and reject if $u\leq\gamma_{N}(T_{N,\alpha},\alpha)$). [^6] For comparison, a non-randomized test of $H_{0}$ with test function $1\{\cdot>T_{N,\alpha}\}$ has CI $\Delta_{(N-T_{N,\alpha})},\infty)$ and rejects only if $M_{0}=0<\Delta_{(N-T_{N,\alpha})}$, implying a more conservative approach. In practice, it suffices to report the interval endpoints along with the rejection probability.
[…]
Figure 3: Rejection probability vs. CI endpoints.
Standard arguments lead to alternative two-sided CIs, which are not, however, dual to the test (2). Observe that $\mathbb{P}_{M}(\Delta_{(i)}\leq M\leq\Delta_{(i+1)})=\pi_{N}(i)$ for $i=0,1,\ldots,N$. Using these probabilities, we can combine intervals $(\Delta_{(i)},\Delta_{(i+1)})$ so as to attain a desired confidence coefficient. Choosing the intervals symmetrically, we obtain intervals
$c_{m,N}=[\Delta_{(1+m)},\Delta_{(N-m)}]$ (9)
satisfying $\mathbb{P}_{M}(M\in c_{m,N})=1-2\Pi_{N}(m)$, where we select $m\in\mathbb{N}$ so that $\Pi_{N}(m)\leq\alpha/2$ for given $\alpha\in(0,1)$. Then, $\mathbb{P}_{M}(M\in c_{m,N})\geq 1-\alpha$, implying a conservative approach. ? extend this argument using randomization, resulting in CIs with exact $(1-\alpha)$ coverage and minimum expected length. Alternative asymptotic CIs can be obtained by bootstrapping (e.g. ? and ?), empirical likelihood (?), sectioning (?) and other methods.
We now develop several useful properties of the population-level feature-importance measure defined by the median $M$. Let $Z=(X,Y)\sim\mathbb{P}$, where $\mathbb{P}$ denotes the population law from which the observations $Z_{i}$ are independently drawn. For concreteness, consider a model that estimates the conditional mean, and let $\mu(X)=\mathbb{E}(Y|X)$ with score function $S$ as given. The (population) feature effect is then $\Delta=S(\mu(\overline{X}),Y)-S(\mu(\underline{X}),Y)$, where $\underline{X}$ and $\overline{X}$ denote appropriate baseline and comparison features, respectively.
As discussed in Section 2.1, the selection of $\underline{X}$ and $\overline{X}$ offers flexibility. Consider the basic choice from Example 2.1,
$\underline{X}=\mathbb{E}(X)\quad\text{and}\quad\overline{X}=(X^{1},\mathbb{E}(X^{2}),\ldots,\mathbb{E}(X^{d})),$ (10)
where, without loss of generality, feature $X^{1}$ is to be tested. Under this specification (10) and given $\mu$ and $S$, the feature effect $\Delta$ is fully determined by the marginal distributions of the features and the joint distribution of the response $Y$ and feature $X^{1}$. Hence, the dependence structure among features does not affect $\Delta$: the importance measure $M$ generated by (10) is robust to feature correlation. This robustness is notable given the well-documented sensitivity of popular importance measures—including SHAP and LOCO—to feature correlation (e.g., ?, ?). Such correlation sensitivity can introduce bias and lead to misleading importance rankings.
Next, we explore the relationship between $M$ and conditional independence. ?, ?, ?, ?, ?, and others develop tests for a conditional-independence hypothesis of feature significance within the model-X knockoff framework of ?. We show that the hypothesis tested by AICO, $M\leq M_{0}=0$, is generally weaker than the alternative conditional-independence hypothesis. Throughout, we use the notation $X^{-1}=(X^{2},\ldots,X^{d})$.
Proposition 4.2.
[…]
A special case where the converse does hold is the standard linear model
$Y=\mu(X)+\epsilon,\quad\text{where}\quad\mu(X)=\beta\cdot X,\quad\beta=(\beta^{1},\ldots,\beta^{d})\in\mathbb{R}^{d}.$ (11)
Before stating the converse of Proposition 4.2 under (11), we note the following “linear agreement” property: under (11) and assuming symmetry of the distribution of $\Delta$, the median $M$ measuring feature importance is proportional to the square of the corresponding regression coefficient. This is precisely the canonical measure of feature importance in a linear model, with proportionality factor given by the feature’s variance.
Proposition 4.3.
Consider the linear model (11). Select (10) and a standard score function $S(m,y)=-(m-y)^{2}$. If $\mathbb{E}(X^{-1}|X^{1})=\mathbb{E}(X^{-1})$ and $\mathbb{E}(\epsilon|X^{1})=0$, we have
$\mathbb{E}(\Delta)=(\beta^{1})^{2}var(X^{1}).$
Proposition $4.$
[…]
5 Synthetic tasks
This section illustrates the robust performance of our test and feature-importance measure on synthetic “known-truth” regression and classification tasks implemented with neural networks. It benchmarks AICO against existing model-agnostic significance tests and variable-selection procedures, as well as against alternative model-specific and model-agnostic measures of feature importance.
5.1 Regression
We consider a feature vector $X=(X^{1},\ldots,X^{d})\subseteq\mathbb{R}^{d}$ of dimension $d=19$ and a response $Y\in\mathbb{R}$ defined by
$Y=\mu(X)+\epsilon,\quad\epsilon\sim N(0,1),$ (12)
where the error $\epsilon$ is independent of $X$, and the conditional mean function $\mu$ is given by
$\mu(x)=3$ $+4x_{1}+x_{1}x_{2}+3x_{3}^{2}+2x_{4}x_{5}+6x_{6}+2\sin(x_{7})$
$+\exp(x_{8})+5x_{9}+3x_{10}+4x_{11}+5x_{12}.$ (13)
The features $X_{1}$ and $X_{6}$ are jointly normal with mean zero, unit variances, and covariance $0.85$. The features $X_{2},\ldots,X_{5}$ and $X_{7}$ are independent $N(0,1)$ variables, and $X_{8}$ is uniformly distributed on [1,1][-1,1]. The binary feature $X_{9}=1\{X_{2}+Z<0\}$ is generated from an independent $Z\sim N(0,1)$. Feature $X_{10}$ follows a Poisson distribution with mean 3, while $X_{11}$ and $X_{12}$ are independent $t$-distributed variables with five degrees of freedom. The non-influential variables $X_{13}$ and $X_{14}$ are independent $N(0,1)$, $(X_{15},X_{16})$ are jointly normal with mean zero, unit variances, and covariance $0.85$, and $X_{17},X_{18}$, and $X_{19}$ are independent $t$-distributed variables with five degrees of freedom.
We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. We train a single-layer neural network model $f$ to estimate the conditional mean $\mu(X)=\mathbb{E}(Y|X)$, defined as
$f(x)=b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x),$ (14)
where $K$ is the number of hidden units, $\psi(y)=1/(1+e^{-y})$ is the activation function, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters to be estimated. Training uses a standard $L_{2}$ loss criterion with $L_{2}$ regularization on the weights $b_{k}$, the Adam optimization method, a batch size of 32, and 500 epochs with early stopping. We iterate across five random initializations. A validation set (25% of $I_{1}$, excluded from training) and a Bayesian search are used to optimize the regularization weight, learning rate, and number of units $K$. The optimal regularization weight is $7\cdot 10^{-4}$, the optimal learning rate is $5\cdot 10^{-4}$, and the optimal number of units is $K=300$. The resulting model achieves an $R^{2}$ exceeding 99% on both the training and test data.
The testing procedure employs the score function $S(m,y)=-(m-y)^{2}$, consistent with the $L_{2}$ training criterion. We use the basic baseline and comparison feature specifications described in Example 2.1. For testing a feature $\ell\in[d]$, we set
$\underline{X}_{i}$ $=(X^{1}_{i},\ldots,X^{\ell-1}_{i},v^{\ell}_{i},X^{\ell+1}_{i},\ldots,X^{d}_{i}),$
$\overline{X}_{i}$ $=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d}),$ (15)
where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ when feature $\ell$ is continuous. For the one-hot feature $\ell=9$, we use $v^{\ell}_{i}=1-X^{\ell}_{i}$, and for the Poisson feature $\ell=10$, we adopt the “alternative mode”
$v^{\ell}_{i}={\arg\max}_{x\in\mathbb{N}\setminus\{X^{\ell}_{i}\}}p^{\ell}(x|I_{1}),$
where $p^{\ell}(\cdot|I_{1})$ denotes the probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. This choice prevents baseline and comparison features from being identical for too many samples, an issue arising with the standard mode. Our specifications ensure that all baseline values lie within the support of $X^{9}$ and $X^{10}$.
$\ell$ AICO CPT HRT LOCO Deep Gaussian
Knockoff Knockoff
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 10 10 10 10
2 10 10 10 10 10 10 10 10 10 10 10 10
3 10 10 10 10 10 10 10 10 10 10 10 10
4 10 10 10 10 10 10 10 10 8 8 10 10
5 10 10 10 10 10 10 10 10 8 8 10 10
6 10 10 10 10 10 10 10 10 10 10 10 10
7 10 10 10 10 10 10 10 10 10 10 10 10
8 10 10 10 10 10 10 10 10 9 9 10 10
9 10 10 10 10 10 10 10 10 10 10 10 10
10 10 10 10 10 10 10 10 10 10 10 10 10
11 10 10 10 10 10 10 10 10 10 10 10 10
12 10 10 10 10 10 10 10 10 10 10 10 10
13 0 0 3 2 0 0 0 0 3 3 4 4
14 0 0 1 1 1 1 1 1 2 2 3 3
15 0 0 2 2 0 0 0 0 0 0 4 4
16 0 0 4 4 1 1 0 0 1 1 3 3
17 0 0 0 0 1 1 1 0 2 2 3 3
18 0 0 0 0 1 1 0 0 0 0 2 2
19 0 0 1 0 0 0 0 0 0 0 2 2
Table 1: Regression model test results for 10 independent training/testing experiments at two standard confidence levels ($\alpha=5\%$ and $\alpha=1\%$). The table compares AICO, CPT, HRT, LOCO, Deep Knockoff, and Gaussian Knockoff procedures, showing the total number of rejections across 10 trials. The test set size is $N=5\cdot 10^{5}$.
We run ten independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report results for confidence levels $\alpha=0.01$ and 0.050.05.777Limited numerical precision can lead to ties, i.e., samples $\Delta_{i}$ whose true values are distinct being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact arises when samples $\Delta_{i}$ with true, nonzero values are rounded to zero. We treat these as nonpositive, excluding them from $n_{+}$—a conservative choice. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. Across all experiments, the test achieves perfect power and produces no false discoveries. Feature correlation (e.g., $\ell=1,6$) and discrete feature distributions (e.g., $\ell=9,10$) do not impair performance. The procedure also correctly identifies the significance of features $\ell=2,4,5$, which influence the response only through interaction terms. Additional experiments confirm robustness to alternative baseline specifications—for example, taking $v_{i}^{\ell}$ as the training mean for the discrete features $\ell=9,10$, or using the median instead of the mean to construct $v_{i}^{\ell}$. We further verify robustness with respect to the score function $S$, including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
[…]
For additional perspective, we also implement strong model-agnostic variable-selection procedures, including the Gaussian Model-X Knockoff method of ? and the Deep Model-X Knockoff method of ?. The Gaussian version assumes a multivariate normal feature distribution, whereas the deep version uses a neural network to learn a generative model of the feature distribution. As with CPT and HRT, the construction of such auxiliary models incurs significant additional computational expense, which AICO avoids.
Table 1 reports results for these alternative procedures across ten independent training/testing experiments. CPT, HRT, and LOCO achieve comparable power to AICO but exhibit higher false discovery rates (some of which, but not all, can be mitigated by a Bonferroni correction). The Gaussian Knockoff procedure consistently selects the significant features 1 through 12 in every experiment, though it produces a small number of false discoveries, consistent with its expected false-discovery-rate control. The Deep Knockoff method reduces false discoveries at the cost of lower power and shows modest difficulty identifying features that enter the model (5.1) through the interaction term $2x_{4}x_{5}$.
We examine the performance of the AICO test as a function of the test-set size $N=|I_{2}|$. Table 2 reports results for $N\in\{10^{5},10^{4},10^{3},500,100\}$; the case $N=5\cdot 10^{5}$ is shown in Table 1. The test maintains full power for sample sizes as small as $N=500$. For smaller $N$, power declines moderately, while false discoveries increase only marginally. These results confirm that the test performs reliably even in small-sample settings.
AICO
$\ell$ $N=10^{5}$ $N=10^{4}$ $N=10^{3}$ $N=500$ $N=250$ $N=100$
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 10 10 10 10
2 10 10 10 10 10 10 10 10 10 9 8 5
3 10 10 10 10 10 10 10 10 10 10 10 10
4 10 10 10 10 10 10 10 10 10 10 9 9
5 10 10 10 10 10 10 10 10 10 10 10 9
6 10 10 10 10 10 10 10 10 10 10 10 10
7 10 10 10 10 10 10 10 10 10 10 10 10
8 10 10 10 10 10 10 10 10 9 8 8 7
9 10 10 10 10 10 10 10 10 10 10 10 10
10 10 10 10 10 10 10 10 10 10 10 10 10
11 10 10 10 10 10 10 10 10 10 10 10 10
12 10 10 10 10 10 10 10 10 10 10 10 10
13 0 0 0 0 0 0 0 0 2 1 0 0
14 0 0 0 0 0 0 1 0 0 0 1 0
15 0 0 1 0 1 1 0 0 1 0 1 0
16 0 0 0 0 0 0 1 1 0 0 1 0
17 1 0 0 0 1 0 0 0 0 0 0 0
18 0 0 1 0 0 0 0 0 0 0 0 0
19 0 0 1 0 0 0 0 0 0 0 2 2
Table 2: Regression model test results for 10 independent training/testing experiments, evaluated at two confidence levels ($\alpha=5\%$ and $1\%$) for several test sample sizes $N$. We report the number of times the null was rejected out of 10 trials for each covariate $\ell\in[d]$.
We now turn to measuring feature importance. Table 3 presents a feature-importance ranking based on the empirical median (6) for one randomly selected training/testing experiment out of ten. It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, along with the standard (non-randomized) two-sided interval (9) and their corresponding coverage probabilities. In addition, Table 3 provides the distribution of the randomized $p$-values. [^9] Ties among samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing confidence intervals and coverage probabilities. This treatment is consistent with the handling of ties in the testing procedure (footnote 7).
AICO, Level $\alpha=1\%$
$p$-value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
Rank Feature Median P(Reject) Lower Upper 98.8% 1.2% Coverage Lower Upper Coverage
1 $X_{9}$ 24.993 1.000 0.000 0.000 24.951 24.951 0.99 24.947 25.040 0.99007
2 $X_{6}$ 15.788 1.000 0.000 0.000 15.661 15.661 0.99 15.647 15.938 0.99007
3 $X_{10}$ 14.572 1.000 0.000 0.000 14.511 14.511 0.99 14.506 14.640 0.99007
4 $X_{12}$ 12.898 1.000 0.000 0.000 12.789 12.789 0.99 12.775 13.022 0.99007
5 $X_{11}$ 8.137 1.000 0.000 0.000 8.063 8.063 0.99 8.054 8.218 0.99007
6 $X_{1}$ 6.210 1.000 0.000 0.000 6.155 6.155 0.99 6.148 6.273 0.99007
7 $X_{3}$ 1.697 1.000 0.000 0.000 1.669 1.669 0.99 1.666 1.730 0.99007
8 $X_{7}$ 0.953 1.000 0.000 0.000 0.941 0.941 0.99 0.940 0.964 0.99007
9 $X_{4}$ 0.416 1.000 0.000 0.000 0.409 0.409 0.99 0.409 0.424 0.99007
10 $X_{5}$ 0.416 1.000 0.000 0.000 0.409 0.409 0.99 0.408 0.424 0.99007
11 $X_{8}$ 0.143 1.000 0.000 0.000 0.140 0.140 0.99 0.139 0.146 0.99007
12 $X_{2}$ 0.084 1.000 0.000 0.000 0.082 0.082 0.99 0.082 0.086 0.99007
- $X_{13}$ -1.154e-07 0.000 0.880 0.881 -2.442e-06 -2.437e-06 0.99 -2.708e-06 1.478e-06 0.99007
- $X_{14}$ -2.054e-07 0.000 0.928 0.928 -2.176e-06 -2.175e-06 0.99 -2.389e-06 9.014e-07 0.99007
- $X_{15}$ -3.060e-07 0.000 0.845 0.846 -3.339e-06 -3.338e-06 0.99 -3.639e-06 2.127e-06 0.99007
- $X_{16}$ -5.512e-07 0.000 0.937 0.937 -2.966e-06 -2.965e-06 0.99 -3.275e-06 9.626e-07 0.99007
- $X_{17}$ -1.340e-06 0.000 0.996 0.997 -3.773e-06 -3.773e-06 0.99 -4.022e-06 0.000 0.99150
- $X_{18}$ 0.000 0.000 0.860 0.861 -1.426e-06 -1.423e-06 0.99 -1.634e-06 1.114e-06 0.99007
- $X_{19}$ 0.000 0.000 0.815 0.816 -1.517e-06 -1.516e-06 0.99 -1.716e-06 1.333e-06 0.99007
Table 3: Regression model feature importance ranking according to the empirical median (6) for one of the independent training/testing experiments that was randomly chosen. The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
The $p$-value distributions show that our approach clearly separates significant from irrelevant features. For the significant features $\ell=1,\ldots,12$, the distributions are concentrated near zero, far below the level $\alpha=1\%$. This provides strong evidence in favor of rejecting the null hypothesis for these features, with rejection probability equal to one. In contrast, the $p$-value distributions of the irrelevant features are concentrated well above $\alpha$, indicating clear retention of the null. No equivocal cases are observed, eliminating the need for randomization.
Feature $\ell=9$ emerges as the most important variable, followed by features $\ell=6$, $10$, and $12$. The remaining significant features are relatively less influential compared with this group. Features $\ell=4$ and $\ell=5$ exhibit (almost) identical medians and confidence intervals, consistent with their symmetric roles in the conditional mean function (5.1).
To demonstrate the stability of AICO feature-importance scores, Figure 4 reports the empirical median (6) for each of ten independent training/testing experiments. The resulting AICO rankings are stable and robust to noise in the data, and remain consistent under alternative choices of the score function $S$, such as S(m,y)=|my|S(m,y)=-|m-y| (results available upon request). For comparison, Figure 4 also displays test-implied feature-importance scores for CPT, HRT, and LOCO (?, ?, ?); kernel and deep SHAP (?); permutation feature-importance scores (?, ?); and Gaussian and deep knockoff importance scores (?, ?).
There is, of course, no ground truth for feature importance (?); we therefore focus on the stability of the rankings rather than their absolute differences across methods. The measures based on formal testing and variable-selection procedures, including AICO, exhibit strong robustness across experiments. In contrast, SHAP delivers inconsistent rankings between its deep and kernel implementations and shows substantial instability across data sets, indicating high sensitivity to noise. These findings raise concerns about the reliability of SHAP-based feature-importance measures.
Figure 4: Feature importance scores across training/testing experiments.
Finally, we compare the computational efficiency of our test with that of the alternative methods discussed above. Figure 5 reports average computation times across ten experiments. [^10] The compute environment consists of 8GB memory and four CPU cores (Intel Xeon CPU E5-2670, 2.60GHz, and Silver 4110 CPU, 2.10GHz). In each experiment we test (AICO, CPT, HRT, LOCO) or screen (knockoffs) all 19 features, or, in the case of SHAP and permutation feature importance, compute importance scores for all 19 features. The reported time includes any training of submodels or auxiliary models, but excludes the time required to train the original neural network (14).
AICO achieves an unprecedented degree of computational efficiency. Among full-fledged testing procedures, it is more than three orders of magnitude faster than LOCO, the next most efficient test. It is nearly two orders of magnitude faster than Gaussian knockoffs, the fastest screening procedure, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method. As shown earlier in this section, AICO’s exceptional computational efficiency does not come at the expense of statistical performance.
Figure 5: Average computation times for testing/screening all 19 features.
5.2 Classification
We next consider a binary response $Y\in\{0,1\}$ governed by the model
$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X)),$ (16)
where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1), and the features $X=(X^{1},\ldots,X^{d})$ are defined as in Section 5.1. We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. The data are imbalanced, as is common in practice (e.g., credit scoring): approximately 10% of the samples satisfy $Y_{i}=1$.
We train a single-layer neural network $f$ to estimate the conditional mean $\mathbb{E}(Y|X)$, specified as
$f(x)=L\!\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right),$ (17)
where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of hidden units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters. Training uses a standard cross-entropy loss with $L_{2}$ regularization (weight $10^{-3}$, approximately optimal in the regression setting), the Adam optimizer, a batch size of 32, and 500 epochs with early stopping. A validation set (25% of $I_{1}$) and a Bayesian search procedure are used to tune the learning rate and number of units $K$. The optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The resulting model achieves an Area Under the ROC Curve exceeding 99% on both the training and test data.
The testing procedure employs the score function $S(m,y)=y\log m+(1-y)\log(1-m)$, consistent with the training criterion. We use the baseline and comparison feature specification (5.1). Table 4 reports the results for confidence levels $\alpha=0.01$ and $0.05$. As in the regression case, the test achieves perfect power and produces no false discoveries. It remains robust to alternative baseline specifications—for example, using the median instead of the mean to construct $v^{\ell}$, or setting $v^{\ell}$ to the training mean even for the one-hot feature $\ell=9$. [^11] The local baseline attains perfect power but yields some false discoveries. The test is also robust to alternative score functions, such as $S(m,y)=-(m-y)^{2}$.
$\ell$ AICO CPT HRT LOCO Deep Gaussian
Knockoff Knockoff
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 9 9 10 10 10 10 10 10
2 10 10 10 10 9 9 10 10 10 10 10 10
3 10 10 10 10 9 9 10 10 10 10 10 10
4 10 10 10 10 9 9 10 10 9 9 10 10
5 10 10 10 10 9 9 10 10 10 10 9 9
6 10 10 10 10 9 9 10 10 10 10 10 10
7 10 10 10 10 9 9 10 10 10 10 10 10
8 10 10 10 10 9 9 10 10 9 9 9 9
9 10 10 10 10 9 9 10 10 10 10 10 10
10 10 10 10 10 9 9 10 10 10 10 6 6
11 10 10 10 10 9 9 10 10 10 10 10 10
12 10 10 10 10 9 9 10 10 10 10 10 10
13 1 1 1 1 2 2 4 4 2 2 1 1
14 0 0 2 1 0 0 5 5 1 1 1 1
15 0 0 2 2 4 4 6 6 0 0 3 3
16 0 0 3 1 2 2 5 5 1 1 2 2
17 1 1 0 0 1 1 5 5 0 0 1 1
18 1 1 0 0 3 3 5 5 0 0 2 2
19 1 1 1 1 1 1 3 3 0 0 1 1
Table 4: Classification model test results for 10 independent training/testing experiments at two standard confidence levels ($\alpha=5\%$ and $\alpha=1\%$). The table compares AICO, LOCO, HRT, Deep Knockoff, and Gaussian Knockoff procedures, showing the total number of rejections across 10 trials. The test set size is $N=5\cdot 10^{5}$.
Table 4 also reports results for LOCO, CPT, HRT, and the Gaussian and deep knockoff procedures. AICO, CPT, and LOCO all achieve full power. LOCO, however, produces a substantial number of false discoveries that cannot be corrected by a standard Bonferroni adjustment. HRT and the knockoff approaches exhibit lower power and some false discoveries. Overall, AICO outperforms all alternatives, matching its performance in the regression setting.
AICO
$\ell$ $N=10^{5}$ $N=10^{4}$ $N=10^{3}$ $N=500$ $N=250$ $N=100$
5% 1% 5% 1% 5% 1% 5% 1% 5% 1% 5% 1%
1 10 10 10 10 10 10 10 10 9 7 3 0
2 10 10 10 10 6 5 3 2 6 5 3 1
3 10 10 10 10 10 10 10 10 10 10 8 8
4 10 10 10 10 10 10 10 9 6 2 4 0
5 10 10 10 10 10 10 8 8 8 6 6 1
6 10 10 10 10 10 10 10 10 8 6 2 1
7 10 10 8 7 3 3 2 0 1 0 0 0
8 10 10 5 2 0 0 0 0 0 0 0 0
9 10 10 10 10 6 3 3 1 5 3 3 1
10 10 10 10 10 10 10 10 10 9 9 8 5
11 10 10 10 10 6 2 9 5 0 0 3 1
12 10 10 10 10 10 9 6 5 5 2 2 0
13 0 0 1 0 1 0 3 0 0 0 0 0
14 0 0 0 0 0 0 1 0 1 0 2 1
15 0 0 0 0 0 0 0 0 1 1 1 0
16 0 0 0 0 0 0 0 0 0 0 2 0
17 1 1 1 0 1 0 2 0 0 0 0 0
18 1 1 1 0 1 0 2 0 0 0 0 0
19 0 0 1 0 1 0 0 0 0 0 0 0
Table 5: Classification model test results for 10 independent training/testing experiments, at two confidence levels ($\alpha=5\%$ and $\alpha=1\%$), across various test sample sizes $N$. We report the total number of rejections over 10 trials.
Table 5 reports test results for varying test-set sample sizes $N$. Power declines more rapidly with smaller $N$ than in the regression case, particularly for features $\ell=7$ and $\ell=8$. The incidence of false discoveries appears largely insensitive to $N$.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 98.8% 1.2% coverage lower upper coverage
1 $X_{3}$ 3.577e-07 1.000 0.000 0.000 3.340e-07 3.340e-07 0.99 3.320e-07 3.829e-07 0.99007
2 $X_{10}$ 2.324e-08 1.000 0.000 0.000 2.077e-08 2.077e-08 0.99 2.053e-08 2.619e-08 0.99007
3 $X_{6}$ 6.722e-10 1.000 0.000 0.000 6.014e-10 6.016e-10 0.99 5.939e-10 7.656e-10 0.99007
4 $X_{1}$ 1.630e-10 1.000 0.000 0.000 1.442e-10 1.443e-10 0.99 1.426e-10 1.862e-10 0.99007
5 $X_{12}$ 6.989e-11 1.000 0.000 0.000 5.826e-11 5.828e-11 0.99 5.702e-11 8.597e-11 0.99007
6 $X_{5}$ 4.339e-11 1.000 0.000 0.000 3.697e-11 3.697e-11 0.99 3.643e-11 5.135e-11 0.99007
7 $X_{4}$ 3.169e-11 1.000 0.000 0.000 2.707e-11 2.707e-11 0.99 2.668e-11 3.733e-11 0.99007
8 $X_{11}$ 1.001e-11 1.000 0.000 0.000 8.027e-12 8.030e-12 0.99 7.821e-12 1.275e-11 0.99007
9 $X_{9}$ 6.787e-12 1.000 0.000 0.000 5.025e-12 5.025e-12 0.99 4.863e-12 9.439e-12 0.99007
10 $X_{2}$ 2.575e-14 1.000 0.000 0.000 1.823e-14 1.825e-14 0.99 1.755e-14 3.704e-14 0.99007
11 $X_{7}$ 1.100e-15 1.000 3.262e-71 3.431e-71 5.257e-16 5.261e-16 0.99 4.808e-16 2.326e-15 0.99007
12 $X_{8}$ 1.344e-17 1.000 1.706e-14 1.743e-14 2.386e-18 2.394e-18 0.99 1.972e-18 4.921e-17 0.99007
- $X_{13}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 1.00000
- $X_{14}$ 0.000 0.000 1.000 1.000 -1.396e-21 -1.390e-21 0.99 -2.680e-21 0.000 0.99485
- $X_{15}$ -3.193e-21 0.000 1.000 1.000 -5.676e-20 -5.672e-20 0.99 -7.258e-20 0.000 0.99503
- $X_{16}$ 0.000 0.000 1.000 1.000 -3.066e-22 -3.052e-22 0.99 -5.372e-22 0.000 0.99501
- $X_{17}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 0.99992
- $X_{18}$ 0.000 0.000 0.940 0.940 0.000 0.000 0.99 -1.849e-32 1.780e-22 0.99007
- $X_{19}$ 0.000 0.000 1.000 1.000 0.000 0.000 0.99 0.000 0.000 0.99999
Table 6: Classification model feature importance ranking according to the empirical median (6) for one of the independent training/testing experiments that was randomly chosen. The table also reports the endpoints of the randomized CI (7) for the median at a $99\%$ level, the standard (non-randomized) two-sided CI (9) along with coverage probabilities, and the distribution of the randomized $p$-value.
Table 6 presents a feature-importance ranking based on the empirical median (6) for one randomly chosen training/testing experiment out of ten. It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, together with the standard (non-randomized) two-sided interval (9) and their coverage probabilities. Table 3 further shows the distribution of the randomized $p$-values.
Figure 6 displays the empirical median (6) across the ten independent training/testing experiments for several methods, including AICO. Across methods, the scores of the less important variables tend to be somewhat more volatile than in the regression setting. As before, SHAP yields inconsistent and highly unstable rankings across experiments.
Figure 6: Classification model feature importance across training/testing experiments.
Figure 7 reports average computation times across ten experiments. As in the regression case, AICO achieves exceptional computational efficiency. It is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is roughly an order of magnitude faster than Gaussian knockoffs, the fastest screening approach, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method.
Figure 7: Average computation times for testing/screening all 19 features.
6 Empirical applications
6.1 Credit card default
This section complements the theoretical and numerical results with an empirical application to credit-card payment data from 30,000 card holders of a Taiwanese issuer, collected in October 2005.121212The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients. The response variable is a binary indicator of payment default, and there are 23 explanatory variables (see Table 7). The dataset is split into training (60%), validation (15%), and testing (25%) subsets, maintaining class proportions through stratified sampling. Categorical variables are one-hot encoded, and continuous variables are standardized to mean 0 and unit variance using training-set statistics.
Feature Data Type Levels (for Categorical) Description
LIMIT_BAL Continuous Amount of the given credit (NT dollars).
SEX Categorical Male, Female Gender of the cardholder.
EDUCATION Categorical Graduate school, University, High school, Others Education level of the cardholder.
MARRIAGE Categorical Married, Single, Others Marital status of the cardholder.
AGE Continuous Age of the cardholder (in years).
[…]
PAY_2 Categorical Same as PAY_0 Repayment status in August 2005.
PAY_3 Categorical Same as PAY_0 Repayment status in July 2005.
PAY_4 Categorical Same as PAY_0 Repayment status in June 2005.
PAY_5 Categorical Same as PAY_0 Repayment status in May 2005.
PAY_6 Categorical Same as PAY_0 Repayment status in April 2005.
BILL_AMT1 Continuous Amount of bill statement (NT dollars) in September 2005.
BILL_AMT2 Continuous Amount of bill statement (NT dollars) in August 2005.
BILL_AMT3 Continuous Amount of bill statement (NT dollars) in July 2005.
BILL_AMT4 Continuous Amount of bill statement (NT dollars) in June 2005.
BILL_AMT5 Continuous Amount of bill statement (NT dollars) in May 2005.
BILL_AMT6 Continuous Amount of bill statement (NT dollars) in April 2005.
PAY_AMT1 Continuous Amount paid (NT dollars) in September 2005.
PAY_AMT2 Continuous Amount paid (NT dollars) in August 2005.
PAY_AMT3 Continuous Amount paid (NT dollars) in July 2005.
PAY_AMT4 Continuous Amount paid (NT dollars) in June 2005.
PAY_AMT5 Continuous Amount paid (NT dollars) in May 2005.
PAY_AMT6 Continuous Amount paid (NT dollars) in April 2005.
Table 7: Raw features of the credit default dataset.
We implement an XGBoost classifier (?) using a cross-entropy loss function. Model performance is evaluated by the area under the ROC curve (AUC) on the validation set, which serves as the criterion for hyper-parameter optimization. We use the Optuna package (?) with the Tree-Structured Parzen Estimator (?), conducting 250 trials. The search space includes the boosting algorithm, tree-specific parameters such as maximum depth and minimum child weight, and regularization terms. Class imbalance is mitigated using a scaling-weight parameter, and training is capped at 1,000 boosting rounds with early stopping after 50 rounds without improvement in validation AUC. The optimal configuration employs gradient tree boosting (?) with an approximation-based tree-building strategy, an $L_{2}$ regularization parameter (lambda) of 0.2, an $L_{1}$ regularization parameter (alpha) of 0.001, a subsampling ratio of 100%, a column-sampling ratio per tree of 65%, a maximum tree depth of six, a minimum child weight of two, a learning rate (eta) of 0.05, a minimum loss reduction (gamma) of 0.00004, and a loss-guided tree-growth policy.
We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the tuned hyper-parameters. Each model is trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC. Ensemble predictions are obtained by averaging the outputs of all 100 models. The resulting ensemble achieves AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set. The results are robust to the model choice; for example, an ensemble of neural networks yields nearly identical outcomes.
As in the classification example of Section 5.2, we use the score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to conduct the test. We continue to apply the baseline and comparison feature specification (5.1), where $v_{i}^{\ell}$ is the mean of $\{X_{k}^{\ell}:k\in I_{1}\}$ when feature $\ell$ is continuous. For categorical variables, we use the alternative mode
$v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1}),$
where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the empirical probability mass function of $\{X_{k}^{\ell}:k\in I_{1}\}$. This specification ensures that the baseline value lies within the categorical support. We treat each categorical variable as a single feature rather than analyzing its associated one-hot encodings separately. If higher resolution is desired, the test can alternatively be run for each individual level of a categorical variable.
Table 8 summarizes the test results for $\alpha=1\%$ (N=7,500N=7{,}500 throughout). The significant features, ranked by the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE. Past delinquencies are strong predictors of future payment default, with more recent payment records exerting greater influence. Actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant. Marital status, gender, and age—but not education level—are significant predictors of default even after controlling for past payment behavior. However, their influence is relatively small, especially for age: the corresponding medians are at least an order of magnitude smaller than those for the payment-history features. These findings suggest that personal circumstances exert a statistically significant, but comparatively modest, influence on borrower behavior beyond patterns in prior payment.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 23.4% 76.6% coverage lower upper coverage
1 PAY_0 0.150 1.000 0.000 0.000 0.143 0.143 0.99 0.142 0.155 0.99063
2 PAY_2 0.121 1.000 0.000 0.000 0.117 0.117 0.99 0.116 0.125 0.99063
3 PAY_3 0.095 1.000 0.000 0.000 0.091 0.091 0.99 0.091 0.099 0.99063
4 PAY_4 0.059 1.000 0.000 0.000 0.055 0.055 0.99 0.054 0.064 0.99063
5 PAY_5 0.037 1.000 2.520e-322 6.423e-322 0.035 0.035 0.99 0.035 0.039 0.99063
6 PAY_6 0.030 1.000 1.020e-297 2.493e-297 0.029 0.029 0.99 0.029 0.032 0.99063
7 MARRIAGE 0.002 1.000 3.912e-11 4.561e-11 0.002 0.002 0.99 0.001 0.003 0.99063
8 SEX 0.002 1.000 1.026e-37 1.383e-37 0.001 0.002 0.99 0.001 0.002 0.99063
9 AGE 3.796e-04 1.000 1.788e-09 2.056e-09 2.200e-04 2.204e-04 0.99 2.036e-04 5.814e-04 0.99063
- LIMIT_BAL 0.000 0.000 0.794 0.800 0.000 0.000 0.99 0.000 3.603e-04 0.99496
- BILL_AMT1 -0.006 0.000 1.000 1.000 -0.007 -0.007 0.99 -0.008 -0.005 0.99063
- BILL_AMT2 -0.001 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -9.540e-04 0.99063
- BILL_AMT3 0.000 0.000 0.975 0.976 -1.599e-04 -1.592e-04 0.99 -1.751e-04 5.542e-05 0.99063
- BILL_AMT4 -4.094e-04 0.000 1.000 1.000 -5.298e-04 -5.284e-04 0.99 -5.504e-04 -2.773e-04 0.99063
- BILL_AMT5 -3.203e-04 0.000 1.000 1.000 -5.070e-04 -5.054e-04 0.99 -5.220e-04 -1.887e-04 0.99063
- BILL_AMT6 -1.176e-04 0.000 1.000 1.000 -2.041e-04 -2.023e-04 0.99 -2.160e-04 -3.890e-05 0.99063
- PAY_AMT1 -0.012 0.000 1.000 1.000 -0.014 -0.014 0.99 -0.014 -0.010 0.99063
- PAY_AMT2 -0.008 0.000 1.000 1.000 -0.009 -0.009 0.99 -0.010 -0.007 0.99063
- PAY_AMT3 -0.006 0.000 1.000 1.000 -0.006 -0.006 0.99 -0.006 -0.005 0.99063
- PAY_AMT4 -3.764e-04 0.000 1.000 1.000 -5.811e-04 -5.779e-04 0.99 -5.998e-04 -1.935e-04 0.99063
- PAY_AMT5 -0.002 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -0.002 0.99063
- PAY_AMT6 -0.001 0.000 1.000 1.000 -0.002 -0.001 0.99 -0.002 -0.001 0.99063
- EDUCATION -3.891e-04 0.000 0.848 0.853 -0.001 -0.001 0.99 -0.001 4.611e-04 0.99063
Table 8: AICO test summary for the credit card default dataset.
We can further explore the role of individual features by running the test on subsets of the test set. For instance, Table 9 reports results for the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$. While EDUCATION is not significant at the $1\%$ level, it becomes significant at the $5\%$ level, as the distribution of its $p$-value is concentrated below $0.05$. When the test is restricted to samples with $\texttt{MARRIAGE}=\text{Married}$ (results available upon request), EDUCATION is no longer significant. These findings indicate that educational level is a significant predictor of default for unmarried borrowers but not for married ones.
AICO, Conditional on MARRIAGE = Single, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper 83.0% 17.0% coverage lower upper coverage
1 PAY_0 0.151 1.000 1.833e-248 5.763e-248 0.143 0.143 0.99 0.142 0.158 0.99023
2 PAY_2 0.130 1.000 0.000 0.000 0.125 0.125 0.99 0.124 0.136 0.99023
3 PAY_3 0.100 1.000 1.072e-295 3.822e-295 0.095 0.095 0.99 0.095 0.105 0.99023
4 PAY_4 0.063 1.000 1.402e-260 4.555e-260 0.056 0.057 0.99 0.056 0.069 0.99023
5 PAY_5 0.040 1.000 4.099e-183 1.071e-182 0.037 0.037 0.99 0.037 0.044 0.99023
6 PAY_6 0.029 1.000 3.112e-171 7.848e-171 0.028 0.028 0.99 0.027 0.031 0.99023
7 AGE 0.002 1.000 2.601e-48 4.147e-48 0.002 0.002 0.99 0.002 0.002 0.99023
8 SEX 0.001 1.000 1.972e-16 2.560e-16 0.001 0.001 0.99 0.001 0.002 0.99023
- LIMIT_BAL -4.292e-04 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 0.000 0.99511
- BILL_AMT1 -0.005 0.000 1.000 1.000 -0.007 -0.007 0.99 -0.007 -0.003 0.99023
- BILL_AMT2 -0.001 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -8.918e-04 0.99023
- BILL_AMT3 4.366e-05 0.000 0.376 0.388 -1.597e-05 -1.038e-05 0.99 -3.777e-05 2.814e-04 0.99023
- BILL_AMT4 -4.339e-04 0.000 1.000 1.000 -6.109e-04 -6.105e-04 0.99 -6.237e-04 -2.882e-04 0.99023
- BILL_AMT5 -1.737e-04 0.000 1.000 1.000 -3.338e-04 -3.327e-04 0.99 -3.612e-04 -3.330e-05 0.99023
- BILL_AMT6 -1.147e-04 0.000 1.000 1.000 -2.326e-04 -2.315e-04 0.99 -2.509e-04 0.000 0.99474
- PAY_AMT1 -0.014 0.000 1.000 1.000 -0.015 -0.015 0.99 -0.016 -0.011 0.99023
- PAY_AMT2 -0.010 0.000 1.000 1.000 -0.012 -0.012 0.99 -0.012 -0.009 0.99023
- PAY_AMT3 -0.006 0.000 1.000 1.000 -0.008 -0.008 0.99 -0.008 -0.005 0.99023
- PAY_AMT4 -4.392e-04 0.000 1.000 1.000 -7.181e-04 -7.118e-04 0.99 -7.488e-04 -1.825e-04 0.99023
- PAY_AMT5 -0.002 0.000 1.000 1.000 -0.003 -0.003 0.99 -0.003 -0.002 0.99023
- PAY_AMT6 -0.002 0.000 1.000 1.000 -0.002 -0.002 0.99 -0.002 -0.001 0.99023
- EDUCATION 0.001 0.000 0.012 0.013 -7.817e-05 -4.987e-05 0.99 -2.034e-04 0.002 0.99023
Table 9: AICO test summary for the credit card default dataset. The test is conditioned the samples for which $\texttt{MARRIAGE}=\text{Single}$.
In the United States, marital status, gender, and age are protected characteristics under Fair Lending laws and therefore are not used—indeed, often not even collected—for estimating default probabilities or credit scores in credit underwriting. Our findings suggest that, within the context of Taiwanese borrowers, these attributes carry predictive information relevant to risk assessment. All else equal, a score constructed without these variables may be less accurate.
6.2 Mortgage risk
We next examine AICO feature significance in a more complex classification setting, predicting the state of a mortgage loan in the next month given a history of time-varying features observed at the beginning of the month. The seven states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned. The analysis is based on an extensive panel dataset of mortgages from four ZIP codes in the greater Los Angeles area, tracked monthly from January 1994 through December 2022. This dataset constitutes a subset of the CoreLogic Loan-Level Market Analytics data. The feature vector has dimension $d=52$ and includes variables such as current loan state, FICO credit score, loan balance, current interest rate, market interest rate, and local unemployment rate, many of which are time-varying.
We apply the data-filtering procedure of ?, resulting in roughly five million loan-month transitions from 117,523 active loans. The training, validation, and test splits are 1/1994–6/2009, 7/2009–12/2009, and 1/2010–12/2022, respectively. We adopt the set–sequence model of ? with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition modeling. Training, validation, and testing procedures follow closely the approach of ?.
The dataset contains missing feature values, which are handled using missingness dummies and forward-filling based on the last observed value for each loan. To avoid biased test results, loan–month samples with missing values are excluded from the test set, as discussed in Appendix C. Because the extent of missingness varies across features, each feature test is conducted on a different effective sample size.
We apply the feature-effect grouping approach described in Appendix B for panel data, treating each mortgage loan as a unit. The panel is unbalanced because loans enter the dataset at different times and are observed for varying durations. The score function $S$ is the negative mean multi-class cross-entropy loss, where the mean is computed across time steps for each loan. This multi-period specification of $S$ enables a joint test of feature importance across all state transitions. Alternative formulations are also possible—for instance, targeting specific transitions such as from Current to Prepaid.
Baseline features are constructed through a two-step aggregation process: first from the loan–month level to the loan level, and then across all loans. For each feature $\ell$ and each loan $k\in I_{1}$ in the training set, the loan-level statistic $u_{k}^{\ell}$ is computed from the loan’s training records $\{X_{k,\cdot}:(k,\cdot)\in I_{1}\}$. Continuous variables are aggregated using the mean, whereas discrete and categorical variables use the mode. Using these loan-level aggregates, baseline and comparison features are then constructed according to (5.1), where $v_{i}^{\ell}$ is the mean of $\{u_{k}^{\ell}:k\in I_{1}\}$ for continuous features. For discrete and categorical variables, we instead use the alternative mode $v_{i}^{\ell}=\arg\max_{u\in\mathcal{X}^{\ell}\setminus\{u_{i}^{\ell}\}}p^{\ell}(u\mid I_{1})$, where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot\mid I_{1})$ is the empirical probability mass function of $\{u_{k}^{\ell}:k\in I_{1}\}$. This two-step aggregation ensures uniform weighting across loans when constructing baseline values, regardless of the number of observations available for each loan.
Table 10 summarizes the test results for $\alpha=1\%$. As expected, the most important feature by far is the current state of the mortgage. The market mortgage rate is the second most important feature, underscoring the influence of time-varying macroeconomic conditions on borrower behavior. Borrowers have a strong incentive to prepay or refinance their mortgage when the market rate falls below their own loan rate, which is ranked nineth in importance. The results also reveal pronounced path dependence: the number of months a borrower remained current and the number of months 30 days delinquent during the past 12 months are both highly significant and economically important predictors of borrower behavior.
AICO, Level $\alpha=1\%$
P-Value Distribution Randomized One-Sided CI Non-Randomized Two-Sided CI
rank variable median P(reject) lower upper lower-1 lower-2 P(lower-2) coverage lower upper coverage missing rate
1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.625 0.99 0.556 0.564 0.9902 0.000
2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.625 0.99 0.005 0.006 0.9902 0.000
3 pool_insurance_flag_value 0.005 1.000 1.031e-240 1.560e-240 0.005 0.005 0.032 0.99 0.005 0.006 0.9903 0.443
4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.625 0.99 0.002 0.003 0.9902 0.000
5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.393 0.99 0.001 0.002 0.9902 0.014
6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.625 0.99 5.230e-04 0.001 0.9902 0.000
7 inferred_collateral_type_value 7.683e-04 1.000 2.425e-70 2.882e-70 6.749e-04 6.750e-04 0.535 0.99 6.652e-04 8.658e-04 0.9902 0.096
8 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.625 0.99 6.662e-04 8.298e-04 0.9903 0.000
9 current_interest_rate_value 6.571e-04 1.000 8.549e-85 1.024e-84 5.789e-04 5.792e-04 0.530 0.99 5.716e-04 7.453e-04 0.9902 0.003
10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.625 0.99 3.341e-04 4.136e-04 0.9902 0.000
11 prepay_penalty_flag_value 2.733e-04 1.000 4.209e-56 4.903e-56 2.365e-04 2.366e-04 0.078 0.99 2.317e-04 3.184e-04 0.9901 0.085
12 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.625 0.99 1.751e-04 2.671e-04 0.9902 0.000
13 buydown_flag_value 2.192e-04 1.000 1.214e-68 1.461e-68 1.882e-04 1.883e-04 0.160 0.99 1.856e-04 2.537e-04 0.9903 0.235
14 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.625 0.99 1.631e-04 2.454e-04 0.9903 0.000
15 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.625 0.99 1.479e-04 2.178e-04 0.9902 0.000
16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.625 0.99 9.399e-05 1.515e-04 0.9902 0.000
17 scheduled_monthly_pi_value 3.616e-05 1.000 4.464e-32 4.979e-32 2.805e-05 2.809e-05 0.067 0.99 2.688e-05 4.472e-05 0.9900 0.003
- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.625 0.99 -4.904e-06 4.245e-05 0.9902 0.000
- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.625 0.99 -4.909e-04 -2.838e-04 0.9902 0.000
- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.625 0.99 -3.507e-05 3.639e-05 0.9902 0.000
- scheduled_principal_value -5.314e-05 0.000 1.000 1.000 -5.989e-05 -5.983e-05 0.980 0.99 -6.061e-05 -4.456e-05 0.9901 0.573
- convertible_flag_value -5.164e-04 0.000 1.000 1.000 -6.553e-04 -6.551e-04 0.034 0.99 -6.722e-04 -3.525e-04 0.9900 0.555
- io_flag_value -4.553e-05 0.000 0.973 0.973 -1.026e-04 -1.020e-04 0.868 0.99 -1.091e-04 1.431e-05 0.9902 0.324
- negative_amortization_flag_value -8.336e-04 0.000 1.000 1.000 -0.001 -0.001 0.825 0.99 -0.001 -4.979e-04 0.9903 0.853
- original_term_value -1.041e-04 0.000 0.999 0.999 -1.843e-04 -1.843e-04 0.139 0.99 -1.924e-04 -1.752e-05 0.9900 1.066e-04
- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.625 0.99 -5.118e-04 -2.708e-04 0.9902 0.000
- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.625 0.99 -5.723e-04 -3.823e-04 0.9902 0.000
- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.625 0.99 -1.846e-04 -9.468e-06 0.9902 0.000
Table 10: AICO test summary for the mortgage risk dataset using the set-sequence model. Training data is grouped and treated at the loan-month level. The missing rate denotes the ratio of feature effects dropped from the test set due to missigness to the total number of loan-level feature effects after grouping (46,890). P(lower-2) represents the probability associated with the larger lower bound (lower-2) of the one-sided randomized CI.
7 Conclusion
This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning. The proposed randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies demonstrate strong statistical performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, illustrating the method’s ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework establishes a foundation for a new standard of interpretability in machine learning—one that combines statistical rigor with computational scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
Appendix A Proofs of technical results
Proof of Proposition 3.2.
Under Assumption 3.1, $n_{+}(x)\sim B_{N,\theta_{x}}$ where $\theta_{x}=1-F_{\Delta}(x)$. The Binomial distribution $B_{N,\theta_{x}}$ has a monotone likelihood ratio, so the Neyman-Pearson lemma (?, Theorem 3.4.1) guarantees the existence of a UMP test for the hypotheses
$\theta_{x}\leq\theta_{0}\quad\text{vs.}\quad\theta_{x}>\theta_{0}$
for some $\theta_{0}\in[0,1]$. Due to the continuity of $F_{\Delta}$, we have $F_{\Delta}(M)=1/2$ so for $\theta_{0}=1/2$ and $x=M_{0}$ these hypotheses are equivalent to (1). The UMP test has test function $\phi(n)=1\{n>T\}+\gamma 1\{n=T\}$ for some $T\in\mathbb{N}$ and $\gamma\in[0,1]$ determined by the requirement that the test’s size be equal to $\alpha$. Formally, for any $M_{0}\in\mathbb{R}$,
$\alpha$ $=\mathbb{E}_{M_{0}}(\phi(n_{+}(M_{0})))$
$=\mathbb{P}_{M_{0}}(n_{+}(M_{0})>T)+\mathbb{E}_{M_{0}}(\gamma|n_{+}(M_{0})=T)\mathbb{P}_{M_{0}}(n_{+}(M_{0})=T).$
It follows that $T=T_{N,\alpha}=q_{1-\alpha}(B_{N,1/2})$ and
$\gamma=\gamma_{N}(T,\alpha)=\frac{\alpha-\mathbb{P}_{M_{0}}(n_{+}(M_{0})>T)}{\mathbb{P}_{M_{0}}(n_{+}(M_{0})=T)}.$
We clearly have $n_{+}(M_{0})\sim B_{N,1/2}$ under $\mathbb{P}_{M_{0}}$ so $\mathbb{P}_{M_{0}}(n_{+}(M_{0})\leq n)=\Pi_{N}(n)$ and furthermore, $\mathbb{P}_{M_{0}}(n_{+}(M_{0})=n)=\pi_{N}(n)$. This gives (3). ∎
Proof of Proposition 4.1.
Let $U\sim U(0,1)$ and $I(x)=1-\phi_{N}(n_{+}(x),\alpha)$, where $n_{+}(x)=\#\{i\in I_{2}:\Delta_{i}>x\}$ for $x\in\mathbb{R}$, as in Proposition 3.2. From ?, a randomized CI for $M$ is given by the $U$-cut of $I(x)$:
$C_{N}(\alpha;U)=\{x\in\mathbb{R}\;|\;I(x)\geq U\}.$
Then, by iterated expectations and the independence of $U$, we get
$\mathbb{P}_{M}(M\in C_{N}(\alpha;U))$ $=\mathbb{P}_{M}(U\leq I(M))$
$=\mathbb{E}_{M}(1-\phi_{N}(n_{+}(M),\alpha))$
$=1-\alpha$
by (4). (Recall that $n_{+}(M)\sim B_{N,1/2}$ under $\mathbb{P}_{M}$.)
Since $n_{+}(M)=N-i$ for $M\in[\Delta_{(i)},\Delta_{(i+1)})$, we have
$C_{N}(\alpha;U)=\{[\Delta_{(i)},\Delta_{(i+1)})\;|\;i=0,\ldots,N;\;1-\phi_{N}(N-i,\alpha)\geq U\}.$
Moreover, note that
$1-\phi_{N}(N-i,\alpha)$ $=1\{N-i\leq T_{N,\alpha}\}-\gamma_{N}(T_{N,\alpha},\alpha)1\{N-i=T_{N,\alpha}\}$
$=1\{N-T_{N,\alpha}<i\}+(1-\gamma_{N}(T_{N,\alpha},\alpha))1\{N-T_{N,\alpha}=i\}.$
Thus $C_{N}(\alpha;U)$ is equal to the union of $\Delta_{(N-T_{N,\alpha})},\Delta_{(N-T_{N,\alpha}+1)})$ and $\Delta_{(N-T_{N,\alpha}+1)},\infty)$ if $1-\gamma_{N}(T_{N,\alpha},\alpha)\geq U$ and $\Delta_{(N-T_{N,\alpha}+1)},\infty)$ otherwise. ∎
Proof of Proposition 4.2.
Conditional independence implies that
$\mu(X)=\mathbb{E}(Y|X)=\mathbb{E}(Y|X^{-1})=g(X^{-1})$
for some function $g$, so $\mu(\cdot)$ does not depend on the first entry of its argument vector. By construction, $\overline{X}$ only differs from $\underline{X}$ at the first entry. Therefore $\mu(\underline{X})=\mu(\overline{X})$ for any pair $(\underline{X},\overline{X})$, and hence the feature effect is identically zero. ∎
Proof of Proposition 4.3.
Letting $\beta^{-1}=(\beta^{2},\ldots,\beta^{d})$, under the stated hypotheses we get
$\mathbb{E}(Y|X^{1})$ $=\beta\cdot\mathbb{E}(X|X^{1})+\mathbb{E}(\epsilon|X^{1})$
$=\beta^{1}X^{1}+\beta^{-1}\cdot\mathbb{E}(X^{-1})$
$=\mu(\overline{X}).$
With the $L_{2}$ score function,
$\Delta$ $=(Y-\mu(\underline{X}))^{2}-(Y-\mu(\overline{X}))^{2}$
$=\left((Y-\mu(\underline{X}))-(Y-\mu(\overline{X}))\right)\left((Y-\mu(\underline{X}))+(Y-\mu(\overline{X}))\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mu(\underline{X})+\mu(\overline{X})-2Y\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mu(\underline{X})-\mu(\overline{X})-2(Y-\mu(\overline{X}))\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}-2\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(Y-\mu(\overline{X})\right).$
Since $\underline{X}$ is deterministic, and $X^{1}$ is the only source of randomness in $\overline{X}$,
$\mathbb{E}(\Delta|X^{1})$ $=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}-2\left(\mu(\underline{X})-\mu(\overline{X})\right)\left(\mathbb{E}(Y|X^{1})-\mu(\overline{X})\right)$
$=\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}.$
By iterated expectations, we conclude that
$\mathbb{E}(\Delta)$ $=\mathbb{E}\left(\left(\mu(\underline{X})-\mu(\overline{X})\right)^{2}\right)$
$=(\beta^{1})^{2}\mathbb{E}\left(\left(\mathbb{E}(X^{1})-X^{1}\right)^{2}\right),$
which completes the proof. ∎
Proof of Proposition $4.$
From Proposition 4.3, we must have that $\mathbb{E}(\Delta)=(\beta^{1})^{2}var(X^{1})=0$, which implies that $\beta^{1}=0$ when $X^{1}$ has strictly positive variance. Therefore,
$Y=\beta^{-1}\cdot X^{-1}+\epsilon.$
The independence of $\epsilon$ and $X$ then implies conditional independence. ∎
Appendix B Panel data
Consider panel data $\{Z_{i,t}\}$, where $i=1,\ldots,m$ indexes units and $t=1,\ldots,T$ indexes time. We partition $\{(i,t):i\in[m],t\in[T]\}$ into two disjoint subsets, $I_{1}$ and $I_{2}$, such that each pair belongs to exactly one subset. This setup accommodates common data-splitting schemes, including temporal splits (distinguishing between in-sample and out-of-sample time periods), unit-based splits (assigning each unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both. For an unbalanced panel, in which units may be observed at different times, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
B.1 Conditioning
Conditioning entails focusing on a suitable subset $I\subseteq I_{2}$ when constructing the feature effect. For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to assess feature significance within a specific borrower subgroup. If some samples in $I_{2}$ are dependent—as is common in panel data—we can restrict attention to the subset $I$ of independent samples $(i,t)\in I_{2}$, albeit at the cost of a smaller test set. We can further partition $I$ by period $t$ to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by time. This yields a sequence of tests and a time series of $p$-values, importance scores, and confidence intervals whose temporal dynamics may provide additional insight.
For constructing baseline and comparison features, the training set $I_{1}$ is treated as a collection of $|I_{1}|$ samples—with unit and time information tied to each—fully exploiting the richness of the panel structure. For each $(i,t)\in I_{2}$, the baseline features can draw information not only from all training samples $I_{1}$ but also from testing samples corresponding to the same unit, $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$. This flexibility enables unit-specific and time-dependent baselines, such as lagged feature values (e.g., the previous month’s feature value) or historical aggregations, thereby enhancing the contextual relevance of the baseline.
B.2 Grouping
An alternative grouping approach enables joint testing across time series. The idea is to aggregate feature effects within a unit of dependence so that dependence is contained within groups. We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$ for $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method. This differs slightly from the conditioning approach: in the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only up to the computation of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, whereas the conditional approach proceeds to construct the feature effect $\Delta_{i,t}$ at the individual-sample level.
Next, we group the computed predictions and corresponding true responses by unit to obtain the unit-level feature effect
$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i}),$
where $S$ is a trajectory-level score function, $\underline{f}_{i}=\{f(\underline{X}_{i,t}):(i,t)\in I_{2}\}$, $\overline{f}_{i}=\{f(\overline{X}_{i,t}):(i,t)\in I_{2}\}$, and $Y_{i}=\{Y_{i,t}:(i,t)\in I_{2}\}$. Unlike the single-period setting, where $S$ evaluates one prediction against one true response, the trajectory-level score function evaluates the entire sequence of predictions against the corresponding sequence of true responses.
Trajectory-level score functions can extend familiar single-period metrics. For instance, one may use the negative mean squared error, averaging across time steps. More complex specifications can exploit richer temporal structure, such as correlation (Information Coefficient, IC), the negative error in maximum drawdown (MDD), S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time-series evaluations tailored to specific domains.
More broadly, the grouping approach extends beyond temporal dependence structures. For spatially dependent data, one may group testing samples by region or neighborhood, under the assumption that neighborhoods are approximately independent. Likewise, for repeated-measures data, grouping by the underlying entity or experimental context can provide an effective way to respect within-group dependence while maintaining valid inference across groups.
B.3 Hybrid approach
Conditioning and grouping can be combined. For instance, suppose we wish to analyze a feature’s impact on model performance over an extended period. A longer period helps smooth short-term noise but makes single-period conditioning insufficient. Conversely, a shorter period preserves data relevance but may preclude grouping. This limitation can be addressed by first conditioning on a time frame of suitable length and then grouping the conditioned data by units corresponding to each time–unit pair. As with conditioning alone, this procedure can be repeated across overlapping time windows, yielding a rolling-window analysis with smoothed single-period results.
Another notable use case of the hybrid method is assessing per-unit consistency of feature performance—how consistent over time is a feature’s impact on model behavior for a given unit? A longer window mitigates short-term fluctuations, and when the multi-period score function aggregates single-period scores, it also captures an additional dimension of per-unit consistency.
For example, consider a multi-period score function defined as the average of a single-period function, such as the negative mean squared error. In this case, the unit-level feature effect $\Delta_{i}$ is effectively the average of the unit–time-level effects $\Delta_{i,t}$ and thus reflects a degree of consistency over time. The longer the time window, the stronger the imposed consistency, as a higher bar is set for variation in feature impact.
Appendix C Missing Data
In practice, many data sets have missing (feature) values. Modelers often impute missing values using a variety of approaches. Sometimes missingness indicator features are added to distinguish imputed from actually observed values in the model training process. In general, we recommend dropping samples with missing/imputed feature values from the test set, assuming these samples can be identified. While this approach reduces the size of the test set and hence the power of the test, it eliminates potentially biased feature effect samples (e.g., feature effects equal to 0, arising when a baseline value is identical to the imputed value such as the mean).
When applying the grouping approach discussed in Appendix B.2, the removal can be performed at the prediction level $f(\cdot)$ instead of the feature effect level $\Delta_{i}$ to preserve sample size. For example, in a panel setting, if a unit $i$ has missing feature values at only one time step tt—i.e., $X^{\ell}_{i,t}$ is missing, but $X^{\ell}_{i,s}$ for $s\neq t$ are present—it may suffice to drop just the predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$ for that specific time step from the group of unit $i$’s predictions. In other words, if only a single point is missing from the time series of feature values, only drop the corresponding point from the time series of predictions, not the entire time series. However, this situation should not be confused with the case when the data is missing in the baseline value but still present in the original data: i.e., $\underline{X}_{i}$ is missing, but $X^{\ell}_{i}$ is present. For example, when using a lagged baseline and data are missing at time t1t-1 but are present at time $t$.
Additionally, a common practice in handling missing data is to add a missingness flag–an additional dummy variable indicating missing feature values. To ensure schematic consistency of the baseline values, the missing flag should always be updated to correspond to current value of the corresponding feature.